Find the amount that $8000 invested at 12% for 6 years will grow to if the interest is compounded.
A) What is it Annually

B) What is it Monthly

Answer by Theo(13342)

8000 is invested for 6 years at 12% per year.

the formula to use is f = p * (1 + r) ^ n

f is the future value
p is the present value
r is the interest rate per time period.
n is the number of time periods.

when the interest rate is compounded annually, the formula becomes:

f = 8000 * (1 + .12) ^ 6 = 15790.58148.

when the interest rate is compounded monthly, the formula becomes:

f = 8000 * (1 + .12/12) ^ (6*12) = 16736.7945.

what is happening is that the annual interest rate of .12 is divided by 12 to get the monthly interest rate of .12/12 and the number of years is multiplied by 12 to get the number of months = 6*12.

the formula uses the rate, not the percent.
the percent divided by 100 equals the rate.